# Chaos controlled and disorder driven phase transitions by breaking permutation symmetry

### v1

@@ Abstract
Introducing disorder in a system typically breaks symmetries and can introduce dramatic changes in its properties such as localization. At the same time, the clean system can have distinct many-body features depending on how chaotic it is. In this work the effect of permutation symmetry breaking by disorder is studied in a system which has a controllable and deterministic regular to chaotic transition. Results indicate a continuous phase transition from an area-law to a volume-law entangled phase irrespective of whether there is chaos or not, as the strength of the disorder is increased. The critical disorder strength obtained by finite size scaling, indicate a strong dependence on whether the clean system is regular or chaotic to begin with. In the process, we also obtain the critical exponents associated with this phase transition. Additionally, we find that a relatively small disorder is seen to be sufficient to delocalize a chaotic system.
@@
Deterministic chaos, originating in the lack of sufficient number of conserved quantities, gives rise to ergodic and mixing dynamics with a positive Lyapunov exponent in the classical case Tabor (1989); Lichtenberg and Lieberman (2013); Ott (2002), and in the quantum case results in states with random matrix properties, thermalization and volume law entanglement Gutzwiller (1991); Stöckmann (1999); Mehta (2004); Wimberger (2014); D’Alessio et al. (2016); Kaufman et al. (2016); Bohigas et al. (1984); Srednicki (1994). Disorder, depending on its relative strength on the other hand, can lead to Griffiths phases Griffiths (1969), Anderson Anderson (1958) or many-body localization Basko et al. (2006); Oganesyan and Huse (2007), and can also result in thermalization and delocalization via breaking symmetries. For example, the Sachdev-Ye-Kitaev (SYK) model is one where randomized all-to-all couplings results in maximal quantum chaos, maximizing a temperature dependent quantum Lyapunov exponent. Sachdev and Ye (1993); Kitaev ; Maldacena and Stanford (2016).
Arrays of superconducting Josephson junction based transmons, which are nonlinear oscillators, are currently the most successful implementations of quantum computers Koch et al. (2007); Barends et al. (2013); Wang et al. (2022). Disorder in the energies of the individual qubit transmons is a necessary evil in the presence of inter-transmon couplings. A detailed study of the consequences for these devices comes with a warning of being “dangerously close to a phase of uncontrollable chaotic fluctuations” Berke et al. (2022).
@@
On the other hand, it has long been appreciated that dynamical chaos has serious effects on practical implementation of quantum many-body systems, as it can increase entanglement to random state values, and small subsystems face effective decoherence, thus questioning the reliability of quantum information processing Zurek (2006); Hauke et al. (2012). Several related aspects have been addressed in Refs. Bandyopadhyay and Lakshminarayan (2002); Gong and Brumer (2003); Flambaum (2000); Song and Shepelyansky (2001); Georgeot and Shepelyansky (2000); Braun (2002); Madhok et al. (2018); Shepelyansky (2001); Piga et al. (2019); Chinni et al. (2021); Berke et al. (2022); Lerose and Pappalardi (2020), for example it has been discussed in Ref. Georgeot and Shepelyansky (2000) that the presence of random individual qubit energies along with residual short range interactions can make the system chaotic under certain limits that compromises the quantum computer. In another work, it is shown that the error in quantum computation may increase exponentially with system size in the chaotic regime in the presence of imperfections Song and Shepelyansky (2001). Thus, it is important to understand the interplay between disorder and dynamical chaos. We therefore, in this work, bring together the clearly connected, yet disparate, fields of chaos and disorder, and study the effect of disorder on regular or chaotic quantum dynamics.
@@
In particular, we investigate the dynamics of $N$ particles originally restricted to the small $O(N)$ permutation symmetric subspace (PSS), but that can have various dynamical regimes, from near-integrable to intermediate and fully chaotic. The disorder breaks the permutation symmetry and results in a phase transition into the full $O(\exp(N))$ Hilbert space. Interestingly, the extent of chaos in the “clean” symmetric space is found to dictate the critical strength of the disorder that ensures the dynamics has penetrated to the full Hilbert space, in the large $N$ limit. The transition has many possible order parameters. The block entanglement changes from area law $\sim\ln N$, characteristic of permutation symmetric states Seshadri et al. (2018) to volume law $\sim N$. Collective observables such as the total angular momentum changes from $\sim N$ to $\sim\sqrt{N}$. This transition shows similarities to a transition from the superradiant to the radiant phases of collective atomic states in a cavity Scully (2015); Gegg et al. (2018); Rubies-Bigorda and Yelin (2022). A recent work studied the dynamics when the system does not start from a permutation symmetric subspace of an otherwise permutation symmetric Hamiltonian Qi et al. (2023). They find non-generic effects of the deep Hilbert space by studying growth of out-of-time order correlator and entanglement. Another work studied the effect of introducing non-uniform initial states on the permutation symmetry of the Hamiltonian Iemini et al. (2024). However, we wish to study the case when the disorder free system starting from PSS has a controlled transition to deterministic chaos, so that its interplay with disorder becomes transparent.
@@
A convenient model for this study is provided by the kicked top which is a textbook example of quantum chaos Haake et al. (1987); D’Ariano et al. (1992); Haake (2013), of one large spin of angular momentum $S=N/2$, that can be considered as the collective dynamics of $N$ spin $1/2$ particles within the permutation symmetric subspace Wang et al. (2004); Madhok et al. (2015). Several experiments have fruitfully implemented this to understand the effects of chaos on entanglement Chaudhury et al. (2009); Neill et al. (2016); Krithika et al. (2019) . A disordered version of the kicked top can be used to investigate the effects of breaking permutation symmetry. The specific Hamiltonian that we study is given by
$H=\frac{k}{\tau 2N}\sum_{\ell<\ell^{\prime}=1}^{N}(1+\epsilon_{\ell\ell^{\prime}})\sigma_{\ell}^{x}\sigma_{\ell^{\prime}}^{x}+\frac{p}{2}\sum_{n=-\infty}^{\infty}\sum_{\ell=1}^{N}\sigma_{\ell}^{y}\delta(t-n\tau),$ (1)
@@
where we have set $\hbar=1$.
@@
The disorder free version $U_{0}$, with all $\epsilon_{\ell\ell^{\prime}}=0$, is the quantum kicked top Haake (2013); Wimberger (2014); Haake et al. (1987); Zyczkowski (1990); Sanders and Milburn (1989); Bhosale and Santhanam (2018); Fox and Elston (1994); Wang et al. (2010); Zou and Wang (2022); Muñoz-Arias et al. (2021); Fiderer and Braun (2018); Bhosale and Santhanam (2017); Sieberer et al. (2019); Lerose and Pappalardi (2020); Kumari and Ghose (2018); Lombardi and Matzkin (2011); Anand et al. (2024) where the Hamiltonian is purely a function of the collective spin operators $J_{\alpha}=\sum_{\ell}^{N}\sigma_{\ell}^{\alpha}/2$ where $\alpha=x,~{}y,~{}z$. The total angular momentum $J^{2}=J_{x}^{2}+J_{y}^{2}+J_{z}^{2}$ defined in terms of these collective spins is a constant of motion in the quantum kicked top Haake (2013); Wang et al. (2004); Ghose et al. (2008). The large $N$ limit of $U_{0}$ is studied as a symplectic classical map on the unit sphere $(X,Y,Z)\mapsto(X^{\prime},Y^{\prime},Z^{\prime})$ where $X=J_{x}/J$, $Y=J_{y}/J$ and $Z=J_{z}/J$. This map shows regular dynamics for small values of $k$, typically $<2$, and fully chaotic dynamics for large values of $k$ ( $k>6$), where the Lyapunov exponent may be approximated as ($\ln k-1$) Constantoudis and Theodorakopoulos (1997); Wang and Robnik (2021). Various aspects of long range interacting systems such as given above are discussed in Ref. Defenu et al. (2024).
@@
If the initial state is an unentangled spin coherent state given by Glauber and Haake (1976); Lee Loh and Kim (2015); Puri (2001)
|$\ket{\psi_{0}}=\ket{\theta,\phi}=\left(\cos\frac{\theta}{2}\ket{0}+e^{i\phi}\sin\frac{\theta}{2}\ket{1}\right)^{\otimes{N}},$ (3)
@@
then the state after $n$ Floquet periods is |$\ket{\psi_{n}}=U_{0}^{n}\ket{\psi_{0}}$. The states at all times $n$ are restricted to the $N+1$ dimensional PSS and conserve the total angular momentum $J^{2}=(N/2)(N/2+1)$. It has also been shown that the chaotic phase can be described using a random matrix theory (RMT) in $N+1$ dimensions Haake (2013). For small values of $N(\leq 4)$, the kicked top is exactly solvable and does yield interesting insights already into the large $N$ cases Ruebeck et al. (2017); Dogra et al. (2019); Sreeram et al. (2021).
@@
Introducing disorder in the system will break the permutation symmetry of the Hamiltonian, and the state $|\psi_{n}\rangle$ after $n$ kicks given by $U_{w}^{n}\ket{\psi_{0}}$ will now span the full $2^{N}$ dimensional Hilbert space where $J^{2}$ is no longer a constant of motion.
Figure 1: Schematic of the dynamics: Top row corresponds to disorder free system where interactions connecting different spins have same magnitude as depicted by connecting arrows of black color. The second term of the Hamiltonian is the periodic kicks along $y-$direction. Bottom row corresponds to a disordered system as shown by connecting arrows of different colors. For times $n\tau<t<(n+1)\tau-\Delta\tau$, the interactions are dominant whereas for $(n+1)\tau-\Delta\tau<t<(n+1)\tau$, it is the field term which is dominant. Here, we use delta kicks which correspond to $\Delta\tau\to 0$ limit. If the radius of the sphere on the right is proportional to $N$, then $\langle J\rangle$ for the disorder free system is restricted on the surface of this sphere, whereas for the disordered system it is pushed inside the sphere where $\langle J\rangle\sim\sqrt{N}$, as also shown in the figure above.
The state |$\ket{\psi_{n}}$ in presence of disorder tends to a random state on the full Hilbert space, denoted as |$\ket{\psi_{\text{RMT}}}$, in which the expectation value of $J_{z}^{2}$ can be estimated by:
@@
where |$\ket{\phi_{j}}=\ket{j_{1}\cdots j_{N}}$, $j_{i}=\pm 1$, are eigenvectors of $J_{z}$ with eigenvalues $\lambda_{j}=(\sum_{\ell}j_{\ell})/2$. The ensemble average over random states $\overline{|\left\langle\psi_{\text{RMT}}|\phi_{j}\right\rangle|^{2}}=1/2^{N}$ is used above along with the fact that $\operatorname{Tr}J_{z}^{2}=N2^{N-1}/2$. Similarly $\left\langle J_{x}^{2}\right\rangle_{\text{RMT}}=\left\langle J_{y}^{2}\right\rangle_{\text{RMT}}=N/4$, and hence $\left\langle J^{2}\right\rangle_{\text{RMT}}=3N/4$. Clearly, ${\langle J^{2}\rangle}$ exhibit a change in the scaling when the dynamics shifts from the permutation symmetric subspace where $\langle J^{2}\rangle\sim N^{2}$ to the full Hilbert space where the state being close to random has a $\sim N$ scaling. It is this transition of the dynamics to the “full” Hilbert space that we shall focus in the current work. A schematic diagram of the Hamiltonian studied highlighting the difference in the dynamics of disorder free as well as the disordered system is shown in Fig. 1.
@@
We have checked numerically that $S_{Q}(A)$ even in the regular regime of small $k$ within permutation symmetric subspace is proportional to ${}\log_{2}(Q+1)$ dependence. A related work calculating the von Neumann entropy in PSS is given in Ref.Castro-Alvaredo and Doyon (2013). On the other hand, when the dynamics traverse the full Hilbert space, $S_{Q}(A)$ is given by the Page value Page (1993); Seshadri et al. (2018); Lee et al. (2017):
@@
As $w$, the standard deviation of the disorder distribution or the disorder strength is increased, various expectation values will potentially shift from their values within the permutation symmetric subspace to the RMT values relevant to the full Hilbert space. This is because increasing the strength of the disorder $w$ will not only take the system out of the permutation symmetric subspace but will also make it more chaotic since increasing $w$ is equivalent to increasing $k$ as can be seen from Eq. 1. Thus as $w\to 0$ or small when the dynamics is predominantly within the permutation symmetric subspace, the scaling of long time limit of various quantities will follow the PSS scaling, $i.e.,$ ${\langle J^{2}\rangle}\sim N^{2}$ and ${\langle{S_{Q}(A)}\rangle}\sim\log_{2}(Q+1)$. In the other extreme limit, when $w\to\infty$ and the dynamics is described by RMT in the full $2^{N}-$dimensions, ${\langle J^{2}\rangle}=3N/4$ and ${\langle{S_{Q}(A)}\rangle}\sim Q$.
We present the results of numerical calculations which are performed with an initial state as given in Eq. 3 with $\theta=2.25,~{}\phi=1.1$ and is evolved using the Floquet operator $U_{w}$ given in Eq. 2 . We use the fast Walsh Hadamard transform to go to system sizes upto $N=16$. The evolution of $\langle J^{2}\rangle$ and $S_{N/2}$ as a function of number of kicks, or time $n$, is shown in Fig. 2 for $k=1$. We find that for small values of $w$, $\langle J^{2}\rangle$ remains a constant till $n\sim N+1$ corresponding to the Heisenberg time relevant to the PSS beyond which the system is taken out of PSS so that its value starts decreasing. It is around $n\sim 2^{N}$, the Heisenberg time relevant to the full Hilbert space that $\langle J^{2}\rangle$ again starts saturating, which implies that the system has realized the presence of full Hilbert space around this time. For all further calculations, we perform time averaging from $n=10^{5}$ to $n=3\times 10^{5}$ when a thermalization has set in for $N=16$, the maximum system size considered in this work. The change in scaling of $J^{2}$ and $S_{N/2}(A)$ as discussed in the text is presented in Fig. 3(a) and (b), respectively, for different $w$.

### v2

# Chaos controlled and disorder driven phase transitions induced by breaking permutation symmetry

@@ Abstract
The effects of disorder and chaos on quantum many-body systems can be superficially similar, yet their interplay has not been sufficiently explored. This work finds a continuous phase transition when disorder breaks permutation symmetry, with details of the transition being controlled by the degree of chaos in the clean limit. The system changes from an area law entangled phase in the permutation symmetric subspace where collective variables exist to volume law entanglement in the full Hilbert space, beyond a critical strength of the disorder. The critical strength tends to zero when the original disorder free system is fully chaotic. We study this mainly via the scaling of the collective spin of non-equilibrium states which transit to have properties of what has been dubbed “deep Hilbert space”. This has potential implications for general many body physics, as well as technologies such as transmon qubits.
I Introduction
@@
Deterministic chaos, originating in the lack of sufficient number of conserved quantities, gives rise to ergodic and mixing dynamics with a positive Lyapunov exponent in the classical case Tabor (1989); Lichtenberg and Lieberman (2013); Ott (2002), and in the quantum case results in states with random matrix properties, thermalization and volume law entanglement Gutzwiller (1991); Stöckmann (1999); Mehta (2004); Wimberger (2014); D’Alessio et al. (2016); Kaufman et al. (2016); Bohigas et al. (1984); Srednicki (1994). Disorder, depending on its relative strength on the other hand, can lead to Griffiths phases Griffiths (1969), Anderson Anderson (1958) or many-body localization Basko et al. (2006); Oganesyan and Huse (2007), and can also result in thermalization and delocalization via breaking symmetries. For example, the extensively investigated Sachdev-Ye-Kitaev (SYK) model, related to the physics of quantum black holes, is one where randomized all-to-all couplings results in maximal quantum chaos Sachdev and Ye (1993); Kitaev ; Maldacena and Stanford (2016). Deterministic chaos can sometimes be modeled as being pseudo-random and insights from disorder physics can be applied as in the well-known connection between dynamical and Anderson localization Grempel et al. (1984).
@@
Arrays of superconducting Josephson junction based transmons, which are nonlinear oscillators, are currently the most successful implementations of quantum computers Koch et al. (2007); Barends et al. (2013); Wang et al. (2022). Disorder in the energies of the individual qubit transmons is a necessary evil in the presence of inter-transmon couplings. A detailed study of the consequences for these devices comes with a warning of being “dangerously close to a phase of uncontrollable chaotic fluctuations” Berke et al. (2022). It has also long been appreciated that dynamical chaos has serious effects on practical implementation of quantum many-body systems, as it can increase entanglement to random state values, and small subsystems face effective decoherence, thus questioning the reliability of quantum information processing Zurek (2006); Hauke et al. (2012); Bandyopadhyay and Lakshminarayan (2002); Gong and Brumer (2003); Flambaum (2000); Song and Shepelyansky (2001); Georgeot and Shepelyansky (2000); Braun (2002); Madhok et al. (2018); Shepelyansky (2001); Piga et al. (2019); Chinni et al. (2021); Lerose and Pappalardi (2020). For example, it has been discussed in Ref. Georgeot and Shepelyansky (2000) that the presence of random individual qubit energies along with residual short range interactions can make the system chaotic under certain limits that compromises the quantum computer loosing its ability to perform well.
Classically it was also found that disorder can “tame” chaos, wherein nonlinear oscillators get synchronized from a chaotic phase in the presence of disorder Braiman et al. (1995). Similarly, many-body localization takes the system from an ergodic phase to a localized one by increasing disorder Lee et al. (2017); Pal and Huse (2010); Oganesyan and Huse (2007); Luitz et al. (2015). It is thus desirable to study the effects of disorder in systems which have a well-established regular $\rightarrow$ chaotic route in the semiclassical or mean-field limit. Here, the transitions in physical properties whose origins lie in dynamical chaos and those from disorder are both at play together.
@@
Towards this end, we investigate the dynamics of $N$ particles originally restricted to the small $O(N)$ permutation symmetric subspace (PSS), but that can have various dynamical regimes, from near-integrable to intermediate and fully chaotic. The disorder breaks the permutation symmetry and results in a phase transition into the full exponentially large, $O(\exp(N))$ Hilbert space. Interestingly, the extent of chaos in the “clean” symmetric space is found to dictate the critical strength of the disorder that ensures the dynamics has penetrated to the full Hilbert space (FHS), in the large $N$ limit. The transition has many possible order parameters. The block entanglement changes from area law $\sim\ln N$, characteristic of permutation symmetric states Seshadri et al. (2018) to volume law $\sim N$. Collective observables such as the total angular momentum changes from $\sim N$ to $\sim\sqrt{N}$. This transition thus shows similarities to a superradiant to radiant phases of collective atomic states in a cavity Scully (2015); Gegg et al. (2018); Rubies-Bigorda and Yelin (2022). Recent studies have explored the effect of introducing non-uniform initial states on the dynamics of a permutation symmetric Hamiltonian Qi et al. (2023); Iemini et al. (2024) which can go into the “deep Hilbert space” Qi et al. (2023), defined as states whose total angular momentum scales as $\sqrt{N}$.
Figure 1: Schematic of the dynamics: Top row corresponds to the disorder free system where interactions connecting different spins have same magnitude $J_{ij}=k$. The second term of the Hamiltonian is the periodic kicks along $y-$direction in uniform magnetic fields. Bottom row corresponds to a disordered system where $J_{ij}$ are random. The angular momentum sphere on the right is a caricature of states with $\langle J\rangle$ for the disorder free system restricted on the surface of this sphere, whereas for the disordered system it is pushed inside the sphere where $\langle J\rangle\sim\sqrt{N}$.
II The model
@@
A convenient model to study the interplay of chaos and disorder is provided by the kicked top which is a textbook example of quantum chaos Haake et al. (1987); D’Ariano et al. (1992); Haake et al. (2018), consisting of one large spin of angular momentum $S=N/2$, that can be considered as the collective dynamics of $N$ spin $1/2$ particles within the permutation symmetric subspace Wang et al. (2004); Madhok et al. (2015). Several experiments have fruitfully implemented this to understand the effects of chaos on entanglement Chaudhury et al. (2009); Neill et al. (2016); Krithika et al. (2019).
A disordered version of the kicked top can be used to investigate the effects of breaking permutation symmetry. The specific Hamiltonian that we study is given by
$H=\frac{k}{2N\tau}\sum_{\ell<\ell^{\prime}=1}^{N}(1+\epsilon_{\ell\ell^{\prime}})\sigma_{\ell}^{x}\sigma_{\ell^{\prime}}^{x}+\frac{p}{2}\sum_{n=-\infty}^{\infty}\sum_{\ell=1}^{N}\sigma_{\ell}^{y}\delta(t-n\tau),$ (1)
@@
where we have set $\hbar=1$. A schematic diagram of the Hamiltonian studied is shown in Fig. 1.
[…]
@@
The disorder free version $U_{0}$, with all $\epsilon_{\ell\ell^{\prime}}=0$, is the quantum kicked top Haake et al. (2018); Wimberger (2014); Haake et al. (1987); Zyczkowski (1990); Sanders and Milburn (1989); Bhosale and Santhanam (2018); Fox and Elston (1994); Wang et al. (2010); Zou and Wang (2022); Muñoz-Arias et al. (2021); Fiderer and Braun (2018); Bhosale and Santhanam (2017); Sieberer et al. (2019); Lerose and Pappalardi (2020); Kumari and Ghose (2018); Lombardi and Matzkin (2011); Anand et al. (2024) where the Hamiltonian is purely a function of the collective spin operators $J_{\alpha}=\sum_{\ell}^{N}\sigma_{\ell}^{\alpha}/2$ where $\alpha=x,~{}y,~{}z$. The total angular momentum $J^{2}=J_{x}^{2}+J_{y}^{2}+J_{z}^{2}$ defined in terms of these collective spins is a constant of motion in the quantum kicked top Haake et al. (2018); Wang et al. (2004); Ghose et al. (2008). Various aspects of disorder free long range interacting systems such as the one given above are discussed in Ref. Defenu et al. (2023, 2024); Pappalardi et al. (2018). The large $N$ limit of $U_{0}$ is studied as a symplectic classical map on the unit sphere $(X,Y,Z)\mapsto(X^{\prime},Y^{\prime},Z^{\prime})$ where $X=J_{x}/J$, $Y=J_{y}/J$ and $Z=J_{z}/J$. This map shows regular dynamics for small values of $k$, typically $<2$, and fully chaotic dynamics for large values of $k$ ( $k>6$), where the Lyapunov exponent may be approximated as ($\ln k-1$) Constantoudis and Theodorakopoulos (1997); Wang and Robnik (2021). The Heisenberg evolution equations of the angular momentum operators are given by:
$\left\langle J_{i}\right\rangle_{n+1}=\left\langle U^{\dagger}J_{i}U\right\rangle_{n}.$ (3)
The classical equations of motion can be obtained in the limit $j\rightarrow\infty$, and are given by :
$\begin{split}X_{n+1}&=Z_{n},\\ Y_{n+1}&=Y_{n}\cos(kZ_{n})+X_{n}\sin(kZ_{n}),\\ Z_{n+1}&=-X_{n}\cos(kZ_{n})+Y_{n}\sin(kZ_{n}).\end{split}$ (4)
The classical phase space map for $k=1$ is shown in Fig. 2, where ($\theta,\phi$) corresponds to the spherical coordinates of the total angular momentum, showing a predominantly regular behavior at this parameter value.
@@
In the quantum case, one can intialize the system in a spin coherent state given by Glauber and Haake (1976); Lee Loh and Kim (2015); Puri (2001)
|$\ket{\psi_{0}}=\ket{\theta,\phi}=\left(\cos\frac{\theta}{2}\ket{0}+e^{i\phi}\sin\frac{\theta}{2}\ket{1}\right)^{\otimes{N}}.$ (5)
@@
It corresponds to a localized state around the classical point $(X,Y,Z)=(\sin\theta\cos\phi,\,\sin\theta\sin\phi,\,\cos\theta)$. The quantum state after $n$ Floquet periods is |$\ket{\psi_{n}}=U_{0}^{n}\ket{\psi_{0}}$. The states at all times $n$ are restricted to the $N+1$ dimensional PSS and conserve the total angular momentum $J^{2}=(N/2)(N/2+1)$. It has also been shown that the chaotic phase can be described using a random matrix theory (RMT) in $N+1$ dimensions Haake et al. (2018). For small values of $N(\leq 4)$, the kicked top is exactly solvable and does yield interesting insights already into the large $N$ cases Ruebeck et al. (2017); Dogra et al. (2019); Sreeram et al. (2021).
@@
Introducing disorder in the system will break the permutation symmetry of the Hamiltonian, and the state $|\psi_{n}\rangle$ after $n$ kicks given by $U_{w}^{n}\ket{\psi_{0}}$ will now be out of PSS where $J^{2}$ is no longer a constant of motion. If the state |$\ket{\psi_{n}}$ in the presence of large disorder tends to a random state on the full Hilbert space, denoted as |$\ket{\psi_{\text{RMT}}}$, the expectation value of $J_{z}^{2}$ can be estimated as:
@@
where |$\ket{\phi_{j}}=\ket{j_{1}\cdots j_{N}}$, $j_{i}=\pm 1$, are eigenvectors of $J_{z}$ with eigenvalues $\lambda_{j}=(\sum_{\ell}j_{\ell})/2$. The ensemble average over random states $\overline{|\left\langle\psi_{\text{RMT}}|\phi_{j}\right\rangle|^{2}}=1/2^{N}$ is used above along with the fact that $\operatorname{Tr}J_{z}^{2}=N2^{N-2}$. Similarly $\left\langle J_{x}^{2}\right\rangle_{\text{RMT}}=\left\langle J_{y}^{2}\right\rangle_{\text{RMT}}=N/4$, and hence $\left\langle J^{2}\right\rangle_{\text{RMT}}=3N/4$. Clearly, ${\langle J^{2}\rangle}$ exhibit a change in the scaling when the dynamics shifts from the permutation symmetric subspace where $\langle J^{2}\rangle\sim N^{2}$ to the FHS where the state being close to random has a $\sim N$ scaling. It is this transition of the dynamics to the “full” Hilbert space that we shall focus in the current work.
@@
We have checked numerically that $S_{Q}(A)$ even in the regular regime of small $k$ within permutation symmetric subspace is proportional to ${}\log_{2}(Q+1)$. A related work calculating the von Neumann entropy in PSS is given in Ref.Castro-Alvaredo and Doyon (2013). On the other hand, when the dynamics traverse the full Hilbert space, $S_{Q}(A)$ is given by the Page value Page (1993); Seshadri et al. (2018); Lee et al. (2017):
@@
Thus, we conclude that as the disorder strength $w$ is increased, various expectation values will potentially shift from their values within the permutation symmetric subspace to the RMT values relevant to the full Hilbert space. This is because increasing the strength of the disorder $w$ will not only take the system out of the permutation symmetric subspace but will also make it more chaotic since increasing $w$ is equivalent to increasing $k$ as can be seen from Eq. 1. Below, we substantiate these change in scalings of $J^{2}$ and $S_{Q}(A)$ using numerics.
III Results and Discussions
It is both instructive and interesting to first examine the evolution of the model in Eq. 1 without the periodic kicks ($p=0$). This is the well-known spin-glass model, the Sherrington-Kirkpatrick model Young (2017); Thirumalai et al. (1989); Pappalardi et al. (2020), with a nonzero average interaction strength. The thermodynamics of this model has long been of great interest. The time evolution of disorder averaged $\langle J^{2}(t)\rangle=\langle\psi_{t}|J^{2}|\psi_{t}\rangle=\langle\psi_{t}|J_{x}^{2}|\psi_{t}\rangle+\langle\psi_{t}|J_{y}^{2}|\psi_{t}\rangle+\langle\psi_{t}|J_{z}^{2}|\psi_{t}\rangle$ for an arbitrary initial coherent state can be solved exactly. Since $J_{x}^{2}$ commutes with the $p=0$ Hamiltonian,
$\langle\psi_{t}|J_{x}^{2}|\psi_{t}\rangle=\langle\theta,\phi|J_{x}^{2}|\theta,\phi\rangle=\frac{N}{4}+\frac{N(N-1)}{4}\cos^{2}{\phi}\sin^{2}{\theta}.$
The calculation of $\langle\psi_{t}|(J_{y}^{2}+J_{z}^{2})|\psi_{t}\rangle$ is more involved. However, a detailed analysis leads to:
$\langle\psi_{t}|(J_{y}^{2}+J_{z}^{2})|\psi_{t}\rangle=\frac{2N}{4}+\frac{N(N-1)}{4}(1-\cos^{2}\phi\sin^{2}\theta)$
[…]
$=\frac{2N}{4}+(1-\cos^{2}\phi\sin^{2}\theta)\exp\left[-w^{2}k^{2}t^{2}(N-2)/N^{2}\right],$
and hence finally
$\begin{split}&\langle J^{2}(t)\rangle=\frac{3N}{4}+\frac{N(N-1)}{4}\left[\cos^{2}\phi\sin^{2}\theta\right.+\\ &\left.(1-\cos^{2}\phi\sin^{2}\theta)\exp\left[-w^{2}k^{2}t^{2}(N-2)/N^{2}\right]\right].\end{split}$ (9)
Figure 3: Evolution of $\left\langle J^{2}(n)\right\rangle$ with time $n$ for the disordered system with and without kicks for two different values of disorder strength $w=1.5,~{}2.6$, initial state |$\ket{\theta,\phi}=\ket{2.25,1.1}$ , $k=1$, and $N=12$. The points correspond to numerical data and the dotted lines correspond to the analytical expression given in Eq. 9 for the unkicked Hamiltonian. The black horizontal lines is $\langle J^{2}_{RMT}\rangle=3N/4=9$ which is well below the saturation value. This saturation value clearly is independent of the strength of disorder, but depends on the initial state, as also predicted by Eq. 9.